Five items was taken from the output of factory. The length and weight of 5 items are given



below.



Length (meter): 5 6 7 9 12



Weight (KG): 13 15 18 19 20



State which of the two characteristics of two items is more variable?


1
Expert's answer
2022-06-10T11:13:12-0400
"mean_x=\mu_x=\dfrac{5+6+7+9+12}{5}=7.8"





"\sigma_x^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{5}(7.84+3.24+0.64""+1.44+17.64)=6.16"





"mean_y=\mu_y=\dfrac{13+15+18+19+20}{5}=17"





"\sigma_y^2=\dfrac{\Sigma(y_i-\bar{y})^2}{n}=\dfrac{1}{5}(16+4+1+4+9)""=6.8"


Since "\sigma_y^2=6.8>6.16=\sigma_x^2," then characteristics of weight is more variable.